Mr. A draws a ticket from a box containing 2 bad, 10 good and 5 very good tickets. If Mr. B draws the next ticket, what is his chance of a better performance than Mr. A?


1
Expert's answer
2021-05-07T09:03:51-0400

There are "2+10+5=17" tickets in a box.


"P(A\ draws \ bad)=\dfrac{2}{17}"

If A draws bad

"P(B\ draws \ good\ or\ very\ good)=\dfrac{15}{16}"


"P(A\ draws \ good)=\dfrac{10}{17}"

If A draws good

"P(B\ draws\ very\ good)=\dfrac{5}{16}"

The probability that Mr. B draws a better ticket


"\dfrac{2}{17}(\dfrac{15}{16})+\dfrac{10}{17}(\dfrac{5}{16})=\dfrac{5}{17}"